OpenStudy (anonymous):

Explain, in complete sentences, which method you would use to solve the following system of equations and why you chose that method. Provide the solution to the system. x - 3y + 2z = -12 x + 2y + 3z = 6 2x - 3y - z = -2

OpenStudy (anonymous):

@teacherman79 @thomaster @TSwizzle @e.mccormick

OpenStudy (anonymous):

which method would you use?

OpenStudy (anonymous):

I don't know, I have been having trouble understanding how to complete these types of questions.

OpenStudy (anonymous):

@teacherman79

OpenStudy (e.mccormick):

The main methods are graphing, elimination, and substitution. This leads to an overview of them all: http://www.purplemath.com/modules/systlin1.htm

OpenStudy (anonymous):

Elimination. Can you help me solve it? @e.mccormick

OpenStudy (e.mccormick):

OK, for elimination you add multiples of one equation to another with the goal of eliminating variables until you have solutions to use.

OpenStudy (e.mccormick):

Lets start by labeling these equations. \(~\bbox[3px,border:2px solid red]{1}~x - 3y + 2z = -12 \) \(~\bbox[3px,border:2px solid red]{2}~x + 2y + 3z = 6 \) \(~\bbox[3px,border:2px solid red]{3}~2x - 3y - z = -2 \) Now, if I take Equation 1 (Eq1) and multiply it by -1 and then add it to Eq2, what do I get?

OpenStudy (e.mccormick):

\(\begin{array} \.-1 \times & x - 3y + 2z = -12 \\ + & x + 2y + 3z = 6 \end{array}\) becomes \(\begin{array} ~& -x + 3y - 2z = 12 \\ + & \;\;\: x + 2y + 3z = 6 \end{array}\) which means: \( x-x + 2y + 3y + 3z - 2z = 6+ 12\) which simplifies to \( 5y + z = 18 \) As you can seem this eliminates the x. That is why they call it elimination. You need to eliminate another x in another equation. Then you can use the two new ones, call them Eq4 and Eq5 to eliminate either a y or a z. Then you will have one of the answers.